OpenStudy (anonymous):

( _______ )^2 = (csc x - 1)(csc x +1) @terenzreignz :) it doesn't have answer choices again! ***my answer: cot(x) or is it tan(x) ?? is that right?

terenzreignz (terenzreignz):

your answer is what now?

OpenStudy (anonymous):

cot(x) ? :/ lol not too sure on this :/

terenzreignz (terenzreignz):

Begin with the Pythagorean identity \[\Large \cos^2(x) +\sin^2(x) = 1\] divide everything by \(\sin^2(x)\) \[\Large \cot^2(x) +1 = \csc^2(x)\] bring 1 to the right side \[\Large \cot^2(x) = \csc^2(x) - 1\] The right side is a difference of two squares... \(x^2-y^2=(x+y)(x-y)\) \[\Large \cot^2(x) = (\csc(x)+1)(\csc(x)-1)\] lawl

OpenStudy (anonymous):

oaky! :P lol :P how do i write that tho into the blank? :/

OpenStudy (anonymous):

cuz it's cot^2(x) right? :/ but the problem says ( ________ )^2 :/

terenzreignz (terenzreignz):

cot of course :P

OpenStudy (anonymous):

ohh and i don't need to write the "x" part? :O

terenzreignz (terenzreignz):

There should already be an x there.

OpenStudy (anonymous):

that's the thing haha... there was no x :/ it just says ( _______ )^2 = (csc x - 1)(csc x +1) :/

OpenStudy (anonymous):

\[\left( \csc x-1 \right)\left( \csc x+1 \right)=\csc ^{2} x-1=\cot ^{2} x= \left( \cot x \right)^{2}\]

OpenStudy (phi):

try cot x

OpenStudy (anonymous):

okay, thanks you guys!! :D